Post Go back to editing

Q. ADPD4101, The power sequence is changed and the green LED generates a peak noise.

Hi Glen.

The circuit was modified so that power is supplied simultaneously according to the power sequence recommended by ADI.

When the green signal is measured in the changed circuit, peak noise occurs every 25 samples.

From the FIFO, 25 samples of each Green/IR/Red are taken every 0.25 seconds, and the peak continues to occur in the first sample.

(51th sample (=25*2+1), 76th sample (=25*3+1), 101, 126, …)

It does not occur in IR and Red.

This problem occurred after modifying the circuit.

FIFO uses I2C communication.

Please refer to the Excel file for details and used registers.

I look forward to your quick response.

Thanks,

Leo

XLSX

Parents
  • Hi Leo,

    Since the spikes showed up only for the green signal, it islikely an issue with the power supply to the green LED. Please check if there are periodic spikes on the VLED line for the green LED.  

    Your setting of FIFO threshold is [0x0006]0x00E0. Change to another value,  i.e. [0x0006]0x0100, and check the signals. If the spikes repeat at a lower rate, then the spikes are associated with FIFO read. Check if there is any interference from the I2C to the VLED.

    You can also switch the order of your timeslot assignment, for instance, put green in timeslot B or C to check if the spikes stay.

    Regards,

    Glen B.

     

  • Hi Glen.

    The customer referred to what you gave us and tested again.

    As a result of testing again, it seems to be a problem with Slot A.

    1. When the customer changes the FIFO threshold to 20 samples, a spike occurs every 20 samples. It seems that the period of the spike depends on the FIFO threshold.

    2. If Slot A is set to IR and Slot B is set to Green, spike noise occurs in IR, which is Slot A, and Green, which is Slot B, operates normally.

    Considering that only Slot A occurs, it is doubtful that there is a problem with I2C communication as well.

    Attach the measurement results.

    I look forward to your quick response.

    Best regards,

    Leo

    XLSX

Reply
  • Hi Glen.

    The customer referred to what you gave us and tested again.

    As a result of testing again, it seems to be a problem with Slot A.

    1. When the customer changes the FIFO threshold to 20 samples, a spike occurs every 20 samples. It seems that the period of the spike depends on the FIFO threshold.

    2. If Slot A is set to IR and Slot B is set to Green, spike noise occurs in IR, which is Slot A, and Green, which is Slot B, operates normally.

    Considering that only Slot A occurs, it is doubtful that there is a problem with I2C communication as well.

    Attach the measurement results.

    I look forward to your quick response.

    Best regards,

    Leo

    XLSX

Children